# exampledefaultouterlinewidth=0pt,innerlinewidth=0pt, roundcorner=5pt,backgroundcolor=myblue, leftmargin=0.2cm,rightmargin=0.2cm

### v1

@@ 7 SGD-AL is an approximate solution
The paper [14] provides a net argument showing that we do not need to work with every $z$ on the contour $\Omega$ defining the integro-differential equation, but only polynomially many in $d$. Recall that $\Omega=\{z\,:\,|z|=\{2\|K\|_{\text{op}},1\}\}$. For a fixed $\xi>0$, we say that ${\Omega}_{\xi}$ is a $d^{-\xi}$-mesh of ${\Omega}$ if ${\Omega}_{\xi}\subset\Omega$ and for every $z\in{\Omega}$ there exists a $\bar{z}\in\Omega_{\xi}$ such that $|z-\bar{z}|<d^{-\xi}$. We can achieve this with $\Omega_{\xi}$ having cardinality, $|\Omega_{\xi}|=C(|\Omega|)d^{\xi}$.
Lemma 7.1 (Net argument, [14], Lemma 5.1).
@@
We also cite the following lemma, which relates two stopping times used throughout this paper.
Lemma 7.2 (Stopping time, [14], Lemma 4.2).
@@
7.1 SGD-AL is an approximate solution
We now show that SGD-AL is an $(d^{\varepsilon},M,T)$-approximate solution.
Proposition 7.1 (SGD-AL is an approximate solution).
@@ 7.1 SGD-AL is an approximate solution
Again, the proof is very similar to [14, Prop. 5.2]. The one difference is that the martingales and error terms are slightly more involved, because of the non-deterministic stepsize we are using. The remainder of this section, along with section 7.2, fills in the details of bounding these lower-order terms, so that the proof can proceed as in [14].
7.1.1 Shorthand notation
@@
The normalization here (dividing by $d$) is chosen so that the $I$ terms are all $O(1)$; this is formally shown in Lemma 7.5.
7.1.2 SGD-AL under the statistic
We follow the approach in [14, Section 5.3] to rewrite the SGD adaptive learning rate update rule as an integral equation. Considering a quadratic function $\varphi:\mathbb{R}^{d}\to\mathbb{R}$ and performing Taylor expansion, we obtain
@@ 7.1.2 SGD-AL under the statistic
From here, we can proceed as in [14, Section 5.3] to show that SGD-AL is an $(\varepsilon,M,T)$-approximated solution.
7.1.3 $S(W_{td},z)$ is an approximate solution
Proof of Proposition 7.1.
@@ 7.1.3 $S(W_{td},z)$ is an approximate solution
(Note that $N$, $I$ and $Q$ can all be found as functions of $S(W_{td},\cdot)$ using contour integration.) It is shown in the proof of [14, Proposition 5.2] that given the analogue of (39) for deterministic stepsize, $S(W_{td},\cdot)$ satisfies
@@
Next, fix a constant $\xi>0$. Let ${\Omega}_{\xi}\subset{\Omega}$ such that there exists a $\bar{z}\in{\Omega}_{\xi}$ such that $|z-\bar{z}|\leq d^{-\xi}$ and the cardinality of ${\Omega}_{\xi}$, $|{\Omega}_{\xi}|=Cd^{\xi}$ where $C>0$ can depend on $\|K\|_{\text{op}}$. For all $z\in{\Omega}$, we note that $\hat{\tau}_{M}\leq\vartheta_{cM}$ (see Lemma 7.2). Consequently, we evaluate the error with the stopped process $W_{td}^{\vartheta}\stackrel{{\scriptstyle\text{def}}}{{=}}W_{d(t\wedge\vartheta)}$ instead of using $\hat{\tau}_{M}$. By Proposition 7.2, the proof of which we have deferred to Section 7.2, we have, for any $\hat{\delta}>0$
@@
An application of the net argument, Lemma 7.1, finishes the proof after setting $\hat{\delta}=\delta/8$ and $\xi=\delta/8$. ∎
7.2 Error bounds
All the martingale and error terms (40) go to 00 as $d$ grows. Formally, we have.
Proposition 7.2.
@@ 7.2 Error bounds
where to suppress notation we use $\vartheta$ as shorthand for $\vartheta_{cM}$, and $c$ is the constant from Lemma 7.2.
@@
This follows from combining Propositions 7.3, 7.4, 7.5, 7.6, and 7.7. ∎
@@
7.2.1 Bounds on the lower order terms in the gradient and hessian
Proposition 7.3 (Hessian error term).
@@ 7.2.1 Bounds on the lower order terms in the gradient and hessian
All the terms inside the expectation, apart from $\|G_{k}\|^{\alpha}_{\infty}$, are deterministic with respect to $\mathcal{F}_{k}$ and bounded by a constant independent of $d$ (see Lemma 7.6). Since we know from Lemma 7.6 that for any $\varepsilon>0$, all moments of $\|G_{k}\|_{\infty}$ are bounded by $d^{\varepsilon}$ w.o.p., we conclude
@@
The last term, ${\mathbb{E}}\,\left[\left|I_{1}(k)\right|^{2}\,|\,\mathcal{F}_{k}\right]^{\frac{1}{2}}$, is also bounded by a constant (see Lemma 7.5), and all together, we find that $|\mathcal{E}_{1}|\leq d^{\varepsilon-\delta/4}$ with overwhelming probability.
@@
The first term is bounded by (6), since $\mathcal{G}_{k}$ and $\operatorname{Tr}(K)I(B_{k})/d$ are bounded independent of $d$; the second term is bounded $Cd^{-1}$ by Lemma 7.9, and the last term is bounded by a constant by Lemma 7.5.
@@
By (6), the first term is bounded by $\hat{C}(\gamma)^{2}(1+\|N_{k}\|_{\infty}^{\alpha}+\|Q_{k}\|_{\infty}^{\alpha}+\|\operatorname{Tr}(K)I(B_{k})/d\|_{\infty}^{\alpha})^{2}$. All of these terms are bounded by a constant independent of $d$ (because of the stopping time.) The second term satisfies the assumptions of Lemma 7.8 with $H=\nabla^{2}\varphi(X_{k})$, and is thus bounded by $Cd^{-1}$. All together,
@@
Proposition 7.4 (Gradient error term).
@@
Just as in the Hessian argument, (6) lets us bound ${\mathbb{E}}\,\left[\left|\gamma(G_{k})+\gamma(\mathcal{G}_{k})\right|^{3}\,|\,\mathcal{F}_{k}\right]^{\frac{1}{4}}$ by $d^{\varepsilon}$ w.o.p., (5) lets us bound ${\mathbb{E}}\,\left[\left|\gamma(G_{k})-\gamma(\mathcal{G}_{k})\right|\,|\,\mathcal{F}_{k}\right]^{\frac{1}{4}}$ by $d^{-\delta/4}$ w.o.p., and Lemma 7.5 lets us bound ${\mathbb{E}}\,\left[\left|I_{3}(k)\right|^{2}\,|\,\mathcal{F}_{k}\right]^{\frac{1}{2}}$ by a constant, giving an overall bound of $|\mathcal{E}_{1}|\leq d^{-\delta/4+\varepsilon}$.
@@
Proposition 7.5 (Gradient martingale).
@@
The second term is superpolynomially small by Lemma 7.5; the first term is superpolynomially small by (6) and (7.6).
@@
Proposition 7.6 (Hessian martingale).
@@
both of which are superpolynomially small by (6) and Lemma 7.5. For the expectation, we have
@@
this product is superpolynomially small by (6), Lemma 7.6, and Lemma 7.5. Overall, we have, with overwhelming probability,
@@
Proposition 7.7 (Integral error term).
@@ 7.2.2 General bounds
Lemma 7.4 (Preliminary bounds).
@@
Proof of (45) in Lemma 7.4.
By [14, Lemma 3.4], if function $f$ is $\alpha$-pseudo-Lipschitz with Lipschitz constant $L(f)$ (as in (2)) and the noise $\epsilon$ is independent of $a$, then
@@
Lemma 7.5 (Gradient and Hessian bounds).
@@
$\leq\frac{C\|W_{k}\|^{2}\|\Delta_{k}\|^{2}}{d}.$ (Lemma 7.3)
Now, $\|W_{k}\|$ is bounded by the stopping time. From Lemma 7.4, $\frac{\|\Delta_{k}\|^{2}}{d}$ is bounded by $d^{\varepsilon}$ w.o.p., and every moment of this expression is bounded independent of $d$, as desired. ∎
Proof of(50) in Lemma 7.5.
@@
$\leq CM^{2}\operatorname{\mathbb{E}}[f^{\prime}(r_{k})^{2}\,|\,\mathcal{F}_{k}].$ (Lemma 7.3)
From Lemma 7.4, $\operatorname{\mathbb{E}}[f^{\prime}(r_{k})^{2}\,|\,\mathcal{F}_{k}]$ is bounded by a constant independent of $d$, as desired. ∎
Proof of (51) in Lemma 7.5.
@@
By Lemma 7.3, $\|\nabla\varphi(X_{k})\|\leq C\|W_{k}\|\leq CM$ (since we are working under a stopping time), and so $\nabla\varphi(X_{k})^{\top}a_{k+1}$ is subgaussian (and thus bounded by $d^{\varepsilon}$ w.o.p.). By (45), $f^{\prime}(r_{k})$ is bounded by $d^{\varepsilon}$ w.o.p., and so their product is bounded by $d^{2\varepsilon}$ w.o.p., as desired. Now for the expectation:
@@
The first term is bounded by a constant independent of $d$, since subgaussian moments are bounded. The second term is bounded by Lemma 7.4, completing the proof. ∎
Lemma 7.6 (Infinity norm bounds).
@@
by Lemma 7.4. ∎
We now prove a concentration result that closely follows [14, Proposition 5.6].
Lemma 7.7 ([14], Lemma 5.2).
@@
To compute the conditional mean $\operatorname{\mathbb{E}}[a_{k+1}\,|\,\hat{\mathcal{F}_{k}}]$ and covariance $(a_{k+1}-\operatorname{\mathbb{E}}[a_{k+1}\,|\,\hat{\mathcal{F}_{k}}])(a_{k+1}-\operatorname{\mathbb{E}}[a_{k+1}\,|\,\hat{\mathcal{F}_{k}}])^{\top}$, we use Lemma 7.7. By Assumption 1, we can write $a_{k+1}=\sqrt{K}v_{k}$, for $v_{k}\sim\mathcal{N}(0,I_{d})$.
@@
$\qquad\leq C\|\mathcal{G}_{k}-\operatorname{Tr}(K)I(B_{k})/d\|_{\infty}$ (Lemma 7.6)
@@
7.3 Specific learning rates
@@
8 Proofs for AdaGrad-Norm analysis
In this section we provide proofs of the propositions related to AdaGrad-Norm in the least squares setting as well as the more general strongly-convex setting. Statements of the propositions for least squares examples are found in Section 4.
8.1 Strongly convex setting
@@
Definition 8.1 ($\hat{L}$-smoothness of outer function $f$).
@@
Definition 8.2 (Restricted Secant Inequality).
@@
Changing variables $u=\Gamma(t)=\int_{0}^{t}\gamma_{s}\mathop{}\!\mathrm{d}s$, we have that $\int_{0}^{\infty}\mathscr{R}(t)\gamma_{t}\mathop{}\!\mathrm{d}t=\int_{0}^{\infty}r(u)\mathop{}\!\mathrm{d}u=\|r\|_{1}$. Rearranging the term in the above equation and taking $t\to\infty$. We obtain: $\|r\|_{1}\leq\frac{\mathscr{D}^{2}(0)}{\left(2\hat{\mu}(f)-(\hat{L}(f))^{2}\frac{1}{d}\operatorname{Tr}(K)\gamma_{0}\right)}$, given that $\frac{\hat{2\mu}(f)}{(\hat{L}(f))^{2}\frac{1}{d}\operatorname{Tr}(K)}>\gamma_{0}$. Using Lemma 8.1, with $i(v)=I(\mathscr{B}(\Gamma^{-1}(v)))={\mathbb{E}}\,_{a,\epsilon}\big{[}(f^{\prime}(x,x^{\star})^{2}\big{]}$ instead of the risk
@@
8.2 Least squares setting

### v2

@@ Appendix C SGD-AL is an approximate solution
The paper [15] provides a net argument showing that we do not need to work with every $z$ on the contour $\Omega$ defining the integro-differential equation, but only polynomially many in $d$. Recall that $\Omega=\{z\,:\,|z|=\max\{2\|K\|_{\text{op}},1\}\}$. For a fixed $\xi>0$, we say that ${\Omega}_{\xi}$ is a $d^{-\xi}$-mesh of ${\Omega}$ if ${\Omega}_{\xi}\subset\Omega$ and for every $z\in{\Omega}$ there exists a $\bar{z}\in\Omega_{\xi}$ such that $|z-\bar{z}|<d^{-\xi}$. We can achieve this with $\Omega_{\xi}$ having cardinality, $|\Omega_{\xi}|=C(|\Omega|)d^{\xi}$.
Lemma C.1 (Net argument, [15], Lemma 5.1).
@@
(We prove in Section C.1 that $S(t,z)$ does indeed satisfy inequality (31).) We also cite the following lemma, which relates two stopping times used throughout this paper.
Lemma C.2 (Stopping time, [15], Lemma 4.2).
@@
C.1 SGD-AL is an approximated solution
Proposition C.1 (SGD-AL is an approximate solution).
@@ C.1 SGD-AL is an approximated solution
Again, the proof is very similar to [15, Prop. 5.2]. The one difference is that the martingales and error terms are slightly more involved, because of the non-deterministic stepsize we are using. The remainder of this section, along with section C.2, fills in the details of bounding these lower-order terms, so that the proof can proceed as in [15].
C.1.1 Shorthand notation
@@
The normalization here (dividing by $d$) is chosen so that the $I$ terms are all $O(1)$; this is formally shown in Lemma C.5.
C.1.2 SGD-AL under the statistic
We follow the approach in [15, Section 5.3] to rewrite the SGD adaptive learning rate update rule as an integral equation. Considering a quadratic function $\varphi:\mathbb{R}^{d}\to\mathbb{R}$ and performing Taylor expansion, we obtain
@@ C.1.2 SGD-AL under the statistic
From here, we can proceed as in [15, Section 5.3] to show that SGD-AL is an $(\varepsilon,M,T)$-approximated solution.
C.1.3 $S(W_{td},z)$ is an approximate solution
Proof of Proposition C.1.
@@ C.1.3 $S(W_{td},z)$ is an approximate solution
(Note that $N$, $I$ and $Q$ can all be found as functions of $S(W_{td},\cdot)$ using contour integration.) It is shown in the proof of [15, Proposition 5.2] that given the analogue of (38) for deterministic stepsize, $S(W_{td},\cdot)$ satisfies
@@
Next, fix a constant $\xi>0$. Let ${\Omega}_{\xi}\subset{\Omega}$ such that there exists a $\bar{z}\in{\Omega}_{\xi}$ such that $|z-\bar{z}|\leq d^{-\xi}$ and the cardinality of ${\Omega}_{\xi}$, $|{\Omega}_{\xi}|=Cd^{\xi}$ where $C>0$ can depend on $\|K\|_{\text{op}}$. For all $z\in{\Omega}$, we note that $\hat{\tau}_{M}\leq\vartheta_{cM}$ (see Lemma C.2). Consequently, we evaluate the error with the stopped process $W_{td}^{\vartheta}\stackrel{{\scriptstyle\text{def}}}{{=}}W_{d(t\wedge\vartheta)}$ instead of using $\hat{\tau}_{M}$. By Proposition C.2, the proof of which we have deferred to Section C.2, we have, for any $\hat{\delta}>0$
@@
An application of the net argument, Lemma C.1, finishes the proof after setting $\hat{\delta}=\delta/8$ and $\xi=\delta/8$. ∎
C.2 Error bounds
All the martingale and error terms (39) go to 00 as $d$ grows. Formally,
Proposition C.2.
@@ C.2 Error bounds
where to suppress notation we use $\vartheta$ as shorthand for $\vartheta_{cM}$, and $c$ is the constant from Lemma C.2.
@@
This follows from combining Propositions C.3, C.4, C.5, C.6, and C.7. ∎
@@
C.2.1 Bounds on the lower order terms in the gradient and hessian
Proposition C.3 (Hessian error term).
@@ C.2.1 Bounds on the lower order terms in the gradient and hessian
All the terms inside the expectation, apart from $\|G_{k}\|^{\alpha}_{\infty}$, are deterministic with respect to $\mathcal{F}_{k}$ and bounded by a constant independent of $d$ (see Lemma C.6). Since we know from Lemma C.6 that for any $\varepsilon>0$, all moments of $\|G_{k}\|_{\infty}$ are bounded by $d^{\varepsilon}$ w.o.p., we conclude
@@
The last term, ${\mathbb{E}}\,\left[\left|I_{1}(k)\right|^{2}\,|\,\mathcal{F}_{k}\right]^{\frac{1}{2}}$, is also bounded by a constant (see Lemma C.5), and all together, we find that $|\mathcal{E}_{1}|\leq d^{\varepsilon-\delta/4}$ with overwhelming probability.
@@
The first term is bounded by (6), since $\mathcal{G}_{k}$ and $\operatorname{Tr}(K)I(B_{k})/d$ are bounded independent of $d$; the second term is bounded $Cd^{-1}$ by Lemma C.9, and the last term is bounded by a constant by Lemma C.5.
@@
By (6), the first term is bounded by $\hat{C}(\gamma)^{2}(1+\|N_{k}\|_{\infty}^{\alpha}+\|Q_{k}\|_{\infty}^{\alpha}+\|\operatorname{Tr}(K)I(B_{k})/d\|_{\infty}^{\alpha})^{2}$. All of these terms are bounded by a constant independent of $d$ (because of the stopping time.) The second term satisfies the assumptions of Lemma C.8 with $H=\nabla^{2}\varphi(X_{k})$, and is thus bounded by $Cd^{-1}$. All together,
@@
Proposition C.4 (Gradient error term).
@@
Just as in the Hessian argument, (6) lets us bound ${\mathbb{E}}\,\left[\left|\gamma(G_{k})+\gamma(\mathcal{G}_{k})\right|^{3}\,|\,\mathcal{F}_{k}\right]^{\frac{1}{4}}$ by $d^{\varepsilon}$ w.o.p., (5) lets us bound ${\mathbb{E}}\,\left[\left|\gamma(G_{k})-\gamma(\mathcal{G}_{k})\right|\,|\,\mathcal{F}_{k}\right]^{\frac{1}{4}}$ by $d^{-\delta/4}$ w.o.p., and Lemma C.5 lets us bound ${\mathbb{E}}\,\left[\left|I_{3}(k)\right|^{2}\,|\,\mathcal{F}_{k}\right]^{\frac{1}{2}}$ by a constant, giving an overall bound of $|\mathcal{E}_{1}|\leq d^{-\delta/4+\varepsilon}$.
@@
Proposition C.5 (Gradient martingale).
@@
The second term is superpolynomially small by Lemma C.5; the first term is superpolynomially small by (6) and (C.6).
@@
Proposition C.6 (Hessian martingale).
@@
both of which are superpolynomially small by (6) and Lemma C.5. For the expectation, we have
@@
this product is superpolynomially small by (6), Lemma C.6, and Lemma C.5. Overall, we have, with overwhelming probability,
@@
Proposition C.7 (Integral error term).
@@ C.2.2 General bounds
Lemma C.4 (Preliminary bounds).
@@
Proof of (44) in Lemma C.4.
By [15, Lemma 3.4], if function $f$ is $\alpha$-pseudo-Lipschitz with Lipschitz constant $L(f)$ (as in (2)) and the noise $\epsilon$ is independent of $a$, then
@@
Lemma C.5 (Gradient and Hessian bounds).
@@
$\leq\frac{C\|W_{k}\|^{2}\|\Delta_{k}\|^{2}}{d}.$ (Lemma C.3)
Now, $\|W_{k}\|$ is bounded by the stopping time. From Lemma C.4, $\frac{\|\Delta_{k}\|^{2}}{d}$ is bounded by $d^{\varepsilon}$ w.o.p., and every moment of this expression is bounded independent of $d$, as desired. ∎
Proof of(49) in Lemma C.5.
@@
$\leq CM^{2}\operatorname{\mathbb{E}}[f^{\prime}(r_{k})^{2}\,|\,\mathcal{F}_{k}].$ (Lemma C.3)
From Lemma C.4, $\operatorname{\mathbb{E}}[f^{\prime}(r_{k})^{2}\,|\,\mathcal{F}_{k}]$ is bounded by a constant independent of $d$, as desired. ∎
Proof of (50) in Lemma C.5.
@@
By Lemma C.3, $\|\nabla\varphi(X_{k})\|\leq C\|W_{k}\|\leq CM$ (since we are working under a stopping time), and so $\nabla\varphi(X_{k})^{\top}a_{k+1}$ is subgaussian (and thus bounded by $d^{\varepsilon}$ w.o.p.). By (44), $f^{\prime}(r_{k})$ is bounded by $d^{\varepsilon}$ w.o.p., and so their product is bounded by $d^{2\varepsilon}$ w.o.p., as desired. Now for the expectation:
@@
The first term is bounded by a constant independent of $d$, since subgaussian moments are bounded. The second term is bounded by Lemma C.4, completing the proof. ∎
Lemma C.6 (Infinity norm bounds).
@@
by Lemma C.4. ∎
We now prove a concentration result that closely follows [15, Proposition 5.6].
Lemma C.7 ([15], Lemma 5.2).
@@
To compute the conditional mean $\operatorname{\mathbb{E}}[a_{k+1}\,|\,\hat{\mathcal{F}_{k}}]$ and covariance $(a_{k+1}-\operatorname{\mathbb{E}}[a_{k+1}\,|\,\hat{\mathcal{F}_{k}}])(a_{k+1}-\operatorname{\mathbb{E}}[a_{k+1}\,|\,\hat{\mathcal{F}_{k}}])^{\top}$, we use Lemma C.7. By Assumption 1, we can write $a_{k+1}=\sqrt{K}v_{k}$, for $v_{k}\sim\mathcal{N}(0,I_{d})$.
@@
$\qquad\leq C\|\mathcal{G}_{k}-\operatorname{Tr}(K)I(B_{k})/d\|_{\infty}$ (Lemma C.6)
@@
C.3 Specific learning rates
@@
Appendix D Proofs for AdaGrad-Norm analysis
In this section we provide proofs of the propositions related to AdaGrad-Norm in the least squares setting as well as the more general strongly convex setting. Statements of the propositions for least squares examples are found in Section 4.
D.1 Strongly convex setting
@@
Definition D.1 ($\hat{L}$-smoothness of outer function $f$).
@@
Definition D.2 (Restricted Secant Inequality).
@@
Changing variables $u=\Gamma(t)=\int_{0}^{t}\gamma_{s}\mathop{}\!\mathrm{d}s$, we have that $\int_{0}^{\infty}\mathscr{R}(t)\gamma_{t}\mathop{}\!\mathrm{d}t=\int_{0}^{\infty}r(u)\mathop{}\!\mathrm{d}u=\|r\|_{1}$. Rearranging the term in the above equation and taking $t\to\infty$. We obtain: $\|r\|_{1}\leq\frac{\mathscr{D}^{2}(0)}{\left(2\hat{\mu}(f)-(\hat{L}(f))^{2}\frac{1}{d}\operatorname{Tr}(K)\gamma_{0}\right)}$, given that $\frac{\hat{2\mu}(f)}{(\hat{L}(f))^{2}\frac{1}{d}\operatorname{Tr}(K)}>\gamma_{0}$. Using Lemma D.1, with $i(v)=I(\mathscr{B}(\Gamma^{-1}(v)))={\mathbb{E}}\,_{a,\epsilon}\big{[}(f^{\prime}(x,x^{\star})^{2}\big{]}$ instead of the risk
@@
D.2 Least squares setting